# Distributed Diamond Formation of Sliding Squares

### v1

@@ Abstract
The sliding square model is a widely used abstraction for studying self-reconfigurable robotic systems, where modules are square-shaped robots that move by sliding or rotating over one another. In this paper, we propose a novel distributed algorithm that allows a group of modules to reconfigure into a diamond shape, starting from an arbitrary side-connected configuration. It is connectivity-preserving and operates under minimal assumptions: one leader module, common chirality, constant memory per module, and visibility and communication restricted to immediate neighbors. Unlike prior work, which relaxes the original sliding square move-set, our approach uses the unmodified move-set, addressing the additional challenge of handling locked configurations. Our algorithm is sequential in nature and operates with a worst-case time complexity of $\mathcal{O}(n^{2})$ rounds, which is optimal for sequential algorithms. To improve runtime, we introduce two parallel variants of the algorithm. Both rely on a spanning tree data structure, allowing modules to make decisions based on local connectivity. Our experimental results show a significant speedup for the first variant, and linear average runtime for the second variant, which is worst-case optimal for parallel algorithms.
@@
Programmable Matter, Distributed Systems, Sliding Squares
@@ 1 Introduction
Edge-connected configurations of square modules, which can reconfigure by sliding over or rotating around one another, are a well-established theoretical model for self-reconfigurable robotic systems. There are numerous applications such as search-and-rescue missions or autonomous construction, where modules collectively navigate complex terrains or form specific shapes while preserving connectivity.
Reconfiguration planning in the sliding square model has primarily focused on centralized approaches. In contrast, we propose a fully distributed algorithm that enables modules to autonomously self-organize into a diamond shape under minimal assumptions about their initial positions, computational capabilities, and sensory range. The diamond serves as an efficient intermediate shape due to its compactness, minimal bounding box, and lack of holes, making it particularly useful for tasks like locomotion, bridge-building, and surface coating.
Our algorithm assumes an initially side-connected configuration with one designated leader module and common chirality among all modules. Modules communicate locally, interact only with adjacent modules, and operate using constant memory. We prove the correctness of our algorithm and analyze its performance in terms of the number of parallel rounds required for reconfiguration.
@@ 1.1 Our Contribution
We introduce a novel distributed reconfiguration algorithm for the sliding square model that transforms any side-connected initial configuration of modules into a diamond shape. The algorithm has a worst-case runtime of $\mathcal{O}(n^{2})$ rounds, which is optimal among solutions that build up any shape sequentially. It operates under strict local communication and visibility constraints, requires only constant memory per module, and maintains connectivity throughout execution. In contrast to previous distributed solutions, which relax the original sliding square move-set, our algorithm utilizes the unmodified move-set, introducing the additional challenge of handling locked configurations. While our main focus is the diamond, the algorithm generalizes to any shape that can be sequentially constructed, provided that at every step, any two cells in the shape remain connected by a shortest path within the shape – a property we formalize as “weak convexity” later in the paper.
@@ 1.2 Related Work
In the limited sliding cube model, in which modules tunnel through the configuration instead of sliding and pivoting, Kawano proposed a series of increasingly efficient reconfiguration algorithms resulting in a linear-time solution [23, 24, 25].
Pivoting models, where modules move only by rotating around corners, introduce significant complexity. Akitaya et al. [1, 6] demonstrated PSPACE-hardness for several pivoting reconfiguration problems in both 2D and 3D. A notable workaround uses five auxiliary modules (called musketeers) to achieve universal reconfiguration in $\mathcal{O}(n^{2})$ pivot moves [4]. Non-universal strategies focus on configurations excluding certain forbidden substructures. Sung et al. [29] identified three such subconfigurations, allowing reconfiguration in $\mathcal{O}(n^{2})$ moves, which was extended by Feshbach and Sung [14] to include non-convex holes.

### v2

# Distributed Rhombus Formation of Sliding Squares

@@ Abstract
The sliding square model is a widely used abstraction for studying self-reconfigurable robotic systems, where modules are square-shaped robots that move by sliding or rotating over one another. In this paper, we propose a novel distributed algorithm that allows a group of modules to reconfigure into a rhombus shape, starting from an arbitrary side-connected configuration. It is connectivity-preserving and operates under minimal assumptions: one leader module, common chirality, constant memory per module, and visibility and communication restricted to immediate neighbors. Unlike prior work, which relaxes the original sliding square move-set, our approach uses the unmodified move-set, addressing the additional challenge of handling locked configurations. Our algorithm is sequential in nature and operates with a worst-case time complexity of $\mathcal{O}(n^{2})$ rounds, which is optimal for sequential algorithms. To improve runtime, we introduce two parallel variants of the algorithm. Both rely on a spanning tree data structure, allowing modules to make decisions based on local connectivity. Our experimental results show a significant speedup for the first variant, and linear average runtime for the second variant, which is worst-case optimal for parallel algorithms.
@@
modular robots, distributed systems, sliding squares
@@ 1 Introduction
Edge-connected configurations of square modules, which can reconfigure by sliding over or rotating around one another (see LABEL:subfig:def-moves), are a well-established theoretical model for self-reconfigurable robotic systems. There are numerous applications such as search-and-rescue missions or autonomous construction, where modules collectively navigate complex terrains or form shapes while preserving connectivity. Reconfiguration planning in the sliding square model has primarily focused on centralized approaches. In contrast, we propose a fully distributed algorithm that enables modules to autonomously self-organize into a rhombus shape under minimal assumptions about their initial positions, computational capabilities, and sensory range. The rhombus serves as an efficient intermediate shape due to its compactness, minimal bounding box, and lack of holes.
@@ 1.1 Our Contribution
We introduce a distributed reconfiguration algorithm for the sliding square model that transforms any initially side-connected configuration with a designated leader and common chirality into a rhombus shape. These two assumptions are minimal but necessary to break initial symmetries. The algorithm has a worst-case runtime of $\mathcal{O}(n^{2})$ rounds, which is optimal among solutions that build up any shape sequentially. It operates under strict local communication and visibility constraints, requires only constant memory per module, and maintains connectivity throughout execution. In contrast to previous distributed solutions, which relax the original sliding square move-set, our algorithm utilizes the unmodified move-set, introducing the additional challenge of handling locked configurations. While our main focus is the rhombus, the algorithm generalizes to any shape that can be sequentially constructed, provided that at every step, any two cells in the shape remain connected by a shortest path within the shape.
@@ 1.2 Related Work
In the limited sliding cube model, modules are restricted to sliding moves only, and reconfiguration is achieved through meta-modules, small groups of modules that act as single units. Kawano proposed a series of increasingly efficient algorithms in this model, culminating in a linear-time solution [23, 24, 25].
Pivoting models, which allow only corner rotations, are significantly harder. Akitaya et al. [1, 6] demonstrated PSPACE-hardness for several pivoting reconfiguration problems in both 2D and 3D. A notable workaround uses five auxiliary modules (called musketeers) to achieve universal reconfiguration in $\mathcal{O}(n^{2})$ pivot moves [4]. Non-universal strategies focus on configurations excluding certain forbidden substructures. Sung et al. [29] identified three such subconfigurations, allowing reconfiguration in $\mathcal{O}(n^{2})$ moves, which was extended by Feshbach and Sung [14] to include non-convex holes.